### v1

@@ 4.3.1 Intersection Systems
Intersection research mainly includes evaluation of the intersection system, vehicle trajectory planning and signal optimization research, as well as the coordinated optimization research of the two. Joerer et al. (2014) introduced a method to evaluate vehicle collision probability at intersections, using modified traffic simulators based on IDM to assess intelligent transportation system safety, especially when traffic rules are disregarded. Schrader and Bittle (2024) employed Sobol global sensitivity analysis to identify how simulation input parameters like IDM’s parameter impact performance metrics of the intersection system. For vehicle trajectory planning related research, IDM is mainly used as a simulation model for background vehicles or manually driven vehicles (Chen et al., 2023b; Dong et al., 2022a; Ghosh and Parisini, 2022; Guo et al., 2023; Jiang et al., 2017; Medina et al., 2018; Stebbins et al., 2017; Wang et al., 2023a; Xiao et al., 2024; Zhang and Du, 2023; Zhang et al., 2023a).In addition,Zhou et al. (2022) used IDM as the microscopic simulation model for CAV under the coorperative intersection control framework. For the signal optimization, IDM is also used as the simulation foundation (Cao et al., 2022a; Sohn and Andert, 2023; Yang and Fan, 2024). Du et al. (2021) developed an efficient coupled vehicle signal control method for optimizing traffic signal timing and CAV trajectories in mixed traffic environments, using IDM as the artificially driven vehicle model. Huang et al. (2023a) proposed a reservation-based cooperative ecological driving model for optimizing CAV trajectories and intersection signals at isolated intersections, using background vehicles simulated by IDM. Li et al. (2018) integrated an electric vehicle eco-driving system and signal control in a V2X environment, using IDM to formulate non-connected electric vehicle following rules and a hybrid algorithm for optimization. Liang et al. (2020) combined connected vehicle technology for traffic signal control at isolated intersections, optimizing platoon dissipation based on IDM estimation of vehicle proximity.
@@ 4.3.2 Ramp Merging
For ramp scenarios, the main focus is on research related to collaborative merging. Baselt et al. (2014); Chen et al. (2023a); Hou et al. (2023); Ito et al. (2019); Ma et al. (2024); Tang et al. (2022b); Xiong et al. (2022); Zhou et al. (2019) proposed various coorperative merging methods by using IDM to simulate a HV in a mixed traffic merging scenario. Gu et al. (2022a) use macroscopic or network fundamental diagrams to study the impact of CAVs on network flow destabilization caused by turning and merging maneuvers. They used IDM, HDM and cooperative IDM to simulate the AV, HV and CAV, respectively. Silgu et al. (2022) proposed a controller integrating ramp metering and variable speed limits for freeway traffic, tested via SUMO simulations with CACC and HVs. For the system evaluation of ramp, Xiao et al. (2018) analyzed the impact of CACC deactivation on traffic flow at merging bottlenecks, using simulations to evaluate capacity and queue discharge rates.
@@ 4.4.1 Vehicle Longitudinal and Lateral control
In individual longitudinal control, scholars tend to design the control algorithm of the automatic vehicle based on physical methods or data-driven methods. Therefore, IDM is still used as a vehicle control model for background vehicles, ACC vehicles, manually driven vehicles or some low-connectivity scenarios (Ju et al., 2023; Kamal et al., 2016, 2022; Li et al., 2022b, 2015; Liu et al., 2023b, d, 2024b; Lu et al., 2018; Ma and Wang, 2022a, b; Manolis et al., 2020; Shi et al., 2021, 2023; Suo et al., 2024; Tang et al., 2022a; Tas et al., 2018; Toghi et al., 2022; Valiente et al., 2024; Wegener et al., 2021; Wu et al., 2022, 2023; Yan et al., 2023; Yang et al., 2016; Yu et al., 2023; Yuan and Yang, 2019; Zhang et al., 2024b, 2023b; Zhou et al., 2020; Milanes and Shladover, 2014; Fernandes and Nunes, 2012; Vivek et al., 2019). Among these, Zhang et al. (2023b) and Manolis et al. (2020) integrated the IDM into the simulation software VISSIM and AIMSUN, respectively.
Compared with longitudinal control, there are relatively few studies on lateral control, and IDM is generally used as a longitudinal control model to study how to perceive, cooperate, make decisions, and adjust in lane change decisions based on the methodologies of game theory, model predictive control, reinforcement learning and etc. like (Chen et al., 2022; Gong et al., 2023; Gu et al., 2022b; Hwang et al., 2022; Li et al., 2024a; Liu et al., 2024a, 2023c; Scheel et al., 2022; Sheng et al., 2023; Sunberg and Kochenderfer, 2022; Wang et al., 2022a; Zhang et al., 2020).
For the individual lateral and longitudinal coordinated control, research usually integrate two types of models into one unified framework (Kanagaraj and Treiber, 2018; Li et al., 2023b; Mullakkal-Babu et al., 2021).
@@ 4.4.2 Evaluation of CAV Systems on Road section
In the road section scale, many research focus on evaluating the performance of the CAV systems. For example,several research evaluate the stability and the capacity of this system (Qin and Wang, 2021; Qin et al., 2023; Sun et al., 2023, 2020; Shang and Stern, 2021; Talebpour et al., 2018; Talebpour and Mahmassani, 2016; Wang et al., 2019; Yao et al., 2021b; Zhou and Zhu, 2021). Alhariqi et al. (2023); Zhou and Zhu (2020) evaluated the impact of different following arrangements in a platoon on the road emissions. Mullakkal-Babu et al. (2022) evaluated the adverse impact of takeover behavior on mixed traffic flow at different takeover time levels for fully automated vehicles. Rahman and Abdel-Aty (2018) explored the impact of CV platooning on highways on longitudinal traffic safety at different market penetration rates, and studied the role of the allocation of CV (i.e., whether they are limited to specific lanes) on traffic safety. Wang et al. (2022d) studied the impact of lane change on the sytem. Wu et al. (2020) investigated the impact of disturbances caused by acceleration and deceleration of the lead vehicle, gusts, and uncertainties in the platoon control system, such as aerodynamic drag and rolling resistance torque, on platoon communication. Yao et al. (2021a) studied the impact of CAVs on fuel consumption and emissions of mixed highway traffic flows. Yao et al. (2023) explored the impact of the maximum size of an autonomous platoon on mixed traffic flows. Zhou et al. (2021) studied the impact of degradation of CAV platooning control modes on the fundamental diagram. Autonomous driving car testing is an important part of autonomous driving research and development, and IDM also plays an important role in this as the background vehicle (Sun et al., 2022a; Li et al., 2022a; Sun et al., 2022b; Cao et al., 2023; Zhang et al., 2024a; Zhang and Sun, 2024).
@@ 4.4.3 Traffic optimization of road section and network
Furthermore, some system management research on the road section and network have been conducted. Some tried to improve the traffic capacity of road section based on platoon control. Most of them use IDM for simulating the HV like Ferreira and d’Orey (2012); Ge and Orosz (2017); He and Wu (2018); Liu et al. (2022b); Montanino and Punzo (2015); Wang et al. (2014, 2022b, 2022c), while other platoon control related research always use IDM as the degradation model of CAV. Some use the VSL to optimize the CAV system in which IDM is used to simulate HV or CAV (Khondaker and Kattan, 2015; Li et al., 2022c, 2017c; Lu et al., 2023; Othman et al., 2022). Mahajan et al. (2024); Schakel and van Arem (2014) proposed a speed-based lane guidance system and a new on-board advisory system, respectively, to provide recommendations on lanes, vehicle speeds, and headway, optimizing lane allocation under high-traffic conditions to reduce capacity degradation. Wang et al. (2020c) studied a cooperative ecological driving system for signal corridors, where the CAV switches between a leading vehicle and a following vehicle, and IDM is used to simulate a CAV with network communication failure.
On the road network scale, various scenarios are simulated with IDM as the underlying model. Huang et al. (2023b) considered different signal control schemes and demand loading patterns, and focused on the overall network level to consider the traffic flow patterns brought about by autonomous driving mixed traffic. Claes et al. (2011) proposed a distributed vehicle path planning method based on predictive information.
@@ 4.5.1 Applications in vehicle communications
In terms of in-vehicle networks, IDM is used as a fundamental model for vehicle operation. Wang et al. (2023b) and Wang et al. (2024b) have significantly advanced the understanding of cyberattacks on automated vehicles by developing an optimal feedback control law and a comprehensive framework for modeling attacks on adaptive cruise control vehicles. Dong et al. (2022b) introduced a resource allocation algorithm for 5G-V2X platoons, optimizing communication performance under finite block length constraints, crucial for ensuring ultra-reliable low-latency communications. Jia et al. (2014a) and Jia et al. (2014b) delved into VANET-enabled platoon architectures, emphasizing disturbance adaptation and network connectivity, which are vital for effective safety message transmission under varying system parameters. Moreover, Segata and Lo Cigno (2013) assessed the integration of automatic emergency braking systems within VANETs, while Tian et al. (2017) proposed a decentralized relay selection algorithm based on non-cooperative game theory to enhance cooperative dynamics.
@@ 4.5.2 Applications in civil engineering
Ge et al. (2022) introduced an advanced traffic flow load monitoring system that integrates machine vision and weigh-in-motion systems, leveraging deep learning for precise vehicle detection and extraction of key parameters for traffic flow load simulation. Zhou et al. (2024) proposed a two-step Hybrid Virtual-Real Traffic Simulation method to replicate the spatiotemporal distribution of bridge loads, demonstrating the broad applicability of IDM in various interdisciplinary fields.
@@ 5.1 Combination of Big Data
Current IDM research faces several limitations. Despite the widespread use of open-source datasets like the NGSIM, Naples data set (Punzo and Simonelli, 2005), Hefei data set (Jiang et al., 2015), and the HighD data set (Krajewski et al., 2018), data scarcity remains a significant challenge, affecting both empirical and experimental data acquisition. The limited availability and scope of these datasets make it difficult to extract universal behavioral or macro-level characteristics. Additionally, insufficient data volume constrains model validation and refinement. This situation not only hampers the development of mechanism-driven models like IDM but also limits the effective training of data-driven models, such as those based on machine learning and deep learning. Issues of data quality and diversity are also prominent, as existing datasets may lack representation of specific traffic conditions or regions. Furthermore, inadequate data collection frequency and precision hinder the ability to capture dynamic traffic changes.
Integrating big data into IDM research can significantly address current limitations through both empirical and experimental data approaches. Empirical data, gathered from traffic monitoring systems, floating car data, and intelligent infrastructure, offers a vast and dynamic dataset. This abundance of data provides a robust foundation for testing and validating models across diverse scenarios. By leveraging these large-scale datasets, IDM can be further refined and integrated with machine learning and deep learning techniques, enhancing traffic prediction capabilities. This integration enables more accurate and adaptive models that can efficiently respond to real-time traffic conditions. On the experimental data front, the advantages are equally compelling. Conducting controlled traffic experiments enables researchers to delve deeper into the mechanisms of driver behavior and traffic flow. Experimental data provides the opportunity to isolate and study specific variables, minimizing interference and uncovering fundamental characteristics and mechanisms (Tian and Jiang, 2023). This detailed insight is crucial for guiding model development and improving theoretical frameworks. By bridging micro-level decision-making processes with macro-level traffic phenomena, experimental data facilitates a comprehensive understanding of traffic dynamics. This approach not only enriches the theoretical foundation of IDM but also supports the development of models that are both innovative and aligned with real-world complexities.
@@ 5.2 Hybrid Modelling Methodology
Hybrid modeling, which combines data-driven and mechanism-driven approaches, offers significant advantages over purely mechanistic or data-driven models. It leverages the strengths of both methodologies: the adaptability and pattern recognition capabilities of data-driven models, and the theoretical rigor and interpretability of mechanism-driven models like (Yuan et al., 2020, 2021). This synergy can lead to more robust and versatile traffic models, capable of addressing complex, real-world scenarios. However, the integration process faces several limitations.
Firstly, the integration often lacks seamless fusion, resulting in limited adaptability and flexibility in practical applications. Secondly, data-driven models and mechanism-driven models often have different underlying structures and assumptions, which makes combining them smoothly a challenge. This can lead to a lack of cohesion in the hybrid model, where components operate in isolation rather than in a unified manner. Additionally, the calibration and tuning processes for each model type can be complex and may not align well, resulting in inefficiencies or conflicts in parameter settings. Additionally, hybrid models can exhibit inconsistent performance across different traffic scenarios, particularly in data-scarce or anomalous situations, affecting their robustness.Moreover, the lack of standardized metrics can lead to unfair comparisons and inconsistent assessments just as Wang et al. (2024a) pointed out. Furthermore, The challenge of integrating heterogeneous data sources persists, as effectively combining different data types to enhance model performance remains complex. Lastly, computational complexity and real-time applicability are significant constraints; integrated models may demand substantial computational resources, hindering their use in real-time traffic management. Addressing these challenges requires developing more flexible integration frameworks and efficient algorithms to support the advancement and application of IDM models.

### v2

@@ 4.3.1 Intersection Systems
Intersection research mainly includes evaluation of the intersection system, vehicle trajectory planning and signal optimization research, as well as the coordinated optimization research of the two. joererVehicularNetworkingPerspectiveEstimating2014a introduced a method to evaluate vehicle collision probability at intersections, using modified traffic simulators based on IDM to assess intelligent transportation system safety, especially when traffic rules are disregarded. schrader_global_2024 employed Sobol global sensitivity analysis to identify how simulation input parameters like IDM’s parameter impact performance metrics of the intersection system. For vehicle trajectory planning related research, IDM is mainly used as a simulation model for background vehicles or manually driven vehicles (chen_human-like_2023; dong_comparative_2022; ghosh_traffic_2022; guoCoordinationConnectedAutomatedVehicles2023a; jiangEcoApproachingIsolatedSignalized2017a; medina_cooperative_2018; stebbinsCharacterisingGreenLightOptimal2017a; wang_learning_2023; xiaoDecisionMakingAutonomousVehiclesRandom2024a; zhangPlatoonCenteredControlEcoDrivingSignalized2023b; zhang_risk-aware_2023).In addition,zhouCooperativeSignalFreeIntersectionControl2022b used IDM as the microscopic simulation model for CAV under the coorperative intersection control framework. For the signal optimization, IDM is also used as the simulation foundation (cao_gain_2022; sohn_waterfilling_2023; yang_enhancing_2024). duCoupledVehicleSignalControlMethod2021b developed an efficient coupled vehicle signal control method for optimizing traffic signal timing and CAV trajectories in mixed traffic environments, using IDM as the artificially driven vehicle model. huang_reservation-based_2023 proposed a reservation-based cooperative ecological driving model for optimizing CAV trajectories and intersection signals at isolated intersections, using background vehicles simulated by IDM. liEcoDrivingSystemElectricVehicles2018a integrated an electric vehicle eco-driving system and signal control in a V2X environment, using IDM to formulate non-connected electric vehicle following rules and a hybrid algorithm for optimization. liang_equitable_2020 combined connected vehicle technology for traffic signal control at isolated intersections, optimizing platoon dissipation based on IDM estimation of vehicle proximity.
@@ 4.3.2 Ramp Merging
For ramp scenarios, the main focus is on research related to collaborative merging. baselt_merging_2014; chen_deep_2023; hou_cooperative_2023; ito_coordination_2019; maProgressiveVirtualRiskBasedVehicle2024a; tangNovelHierarchicalCooperativeMerging2022b; xiongManagingMergingCAVLane2022b; zhou_state-constrained_2019 proposed various coorperative merging methods by using IDM to simulate a HV in a mixed traffic merging scenario. guNetworkTrafficInstabilityAutomated2022b use macroscopic or network fundamental diagrams to study the impact of CAVs on network flow destabilization caused by turning and merging maneuvers. They used IDM, HDM and cooperative IDM to simulate the AV, HV and CAV, respectively. silgu_combined_2022 proposed a controller integrating ramp metering and variable speed limits for freeway traffic, tested via SUMO simulations with CACC and HVs. For the system evaluation of ramp, xiaoUnravellingEffectsCooperativeAdaptive2018a analyzed the impact of CACC deactivation on traffic flow at merging bottlenecks, using simulations to evaluate capacity and queue discharge rates.
@@ 4.4.1 Vehicle Longitudinal and Lateral control
In individual longitudinal control, scholars tend to design the control algorithm of the automatic vehicle based on physical methods or data-driven methods. Therefore, IDM is still used as a vehicle control model for background vehicles, ACC vehicles, manually driven vehicles or some low-connectivity scenarios (juPredictiveCruiseControllerElectric2023a; kamal_efficient_2016; kamalLookAheadDrivingSchemesEfficient2022a; li_deep_2022; li_effect_2015; liu_semantic_2023; liu_potential_2023; liuSpatiotemporalTrajectoryPlanningAutonomous2024a; luLearningDriverSpecificBehaviorOvertaking2018a; ma_energetic_2022; maPersonalizedDrivingBehaviorsFuel2022a; manolis_real_2020; shiConnectedAutomatedVehicleCooperative2021b; shiDeepReinforcementLearningBased2023b; suo_model-free_2024; tangHighwayDecisionMakingMotionPlanning2022a; tas_making_2018; toghi_social_2022; valiente_prediction-aware_2024; wegener_automated_2021; wuFlowModularLearningFramework2022a; wu_hybrid_2023; yan_unified_2023; yangIsolatedIntersectionControlVarious2016a; yuAssessmentDynamicPlatoonInformation2023a; yuanPredictiveEnergyManagementStrategy2019; zhang_stackelberg_2024; zhangNoMoreRoadBullying2023a; zhou_development_2020; milanesModelingCooperativeAutonomousAdaptive2014b; fernandes_platooning_2012; vivekCyberphysicalRisksHackedInternetconnected2019a). Among these, zhangNoMoreRoadBullying2023a and manolis_real_2020 integrated the IDM into the simulation software VISSIM and AIMSUN, respectively.
Compared with longitudinal control, there are relatively few studies on lateral control, and IDM is generally used as a longitudinal control model to study how to perceive, cooperate, make decisions, and adjust in lane change decisions based on the methodologies of game theory, model predictive control, reinforcement learning and etc. like (chen_adversarial_2022; gongGameTheoryBasedDecisionMakingIterative2023a; guIntegratedEcoDrivingAutomationIntelligent2022a; hwang_autonomous_2022; li_bounded_2024; liuMultistepCooperativeLaneChange2024a; liu_interaction-aware_2023; scheel_recurrent_2022; sheng_cooperation-aware_2023; sunberg_improving_2022; wangModelingBoundedRationalityDiscretionary2022a; zhang_game_2020).
For the individual lateral and longitudinal coordinated control, research usually integrate two types of models into one unified framework (kanagaraj_self-driven_2018; li_developing_2023; mullakkal-babu_hybrid_2021).
@@ 4.4.2 Evaluation of CAV Systems on Road section
In the road section scale, many research focus on evaluating the performance of the CAV systems. For example,several research evaluate the stability and the capacity of this system (qin_analytical_2021; qinStabilityAnalysisConnectedVehicles2023a; sunStabilityExtensionCarFollowingModel2023b; sunRelationshipCarFollowingString2020b; shang_impacts_2021; talebpourEffectInformationAvailabilityStability2018a; talebpourInfluenceConnectedAutonomousVehicles2016a; wang_stability_2019; yao_linear_2021; zhou_analytical_2021). alhariqiImpactVehicleArrangementMixed2023a; zhou_modeling_2020 evaluated the impact of different following arrangements in a platoon on the road emissions. mullakkal-babu_comparative_2022 evaluated the adverse impact of takeover behavior on mixed traffic flow at different takeover time levels for fully automated vehicles. rahman_longitudinal_2018 explored the impact of CV platooning on highways on longitudinal traffic safety at different market penetration rates, and studied the role of the allocation of CV (i.e., whether they are limited to specific lanes) on traffic safety. wangEgoEfficientLaneChangesConnected2022a studied the impact of lane change on the sytem. wuTimeDependentPerformanceAnalysis80211pBased2020a investigated the impact of disturbances caused by acceleration and deceleration of the lead vehicle, gusts, and uncertainties in the platoon control system, such as aerodynamic drag and rolling resistance torque, on platoon communication. yao_fuel_2021 studied the impact of CAVs on fuel consumption and emissions of mixed highway traffic flows. yaoAnalysisImpactMaximumPlatoon2023a explored the impact of the maximum size of an autonomous platoon on mixed traffic flows. zhou_impact_2021 studied the impact of degradation of CAV platooning control modes on the fundamental diagram. Autonomous driving car testing is an important part of autonomous driving research and development, and IDM also plays an important role in this as the background vehicle (sun_scenario-based_2022; li_scegene_2022; sun_adaptive_2022; cao_autonomous_2023; zhang_accelerated_2024; zhang_bayesian_2024).
@@ 4.4.3 Traffic optimization of road section and network
Furthermore, some system management research on the road section and network have been conducted. Some tried to improve the traffic capacity of road section based on platoon control. Most of them use IDM for simulating the HV like ferreira_impact_2012; ge_optimal_2017; heEcoDrivingAdvisoryStrategiesPlatoon2018a; liu_joint_2022; montaninoTrajectoryDataReconstructionSimulationBased2015b; wangRollingHorizonControlFramework2014b; wang_optimal_2022; wang_trajectory_2022, while other platoon control related research always use IDM as the degradation model of CAV. Some use the VSL to optimize the CAV system in which IDM is used to simulate HV or CAV (khondakerVariableSpeedLimitMicroscopic2015a; li_developing_2022; li_integrated_2017; luTD3LVSLLaneLevelVariableSpeed2023a; othman_analysis_2022). mahajan_improving_2024; schakel_improving_2014 proposed a speed-based lane guidance system and a new on-board advisory system, respectively, to provide recommendations on lanes, vehicle speeds, and headway, optimizing lane allocation under high-traffic conditions to reduce capacity degradation. wang_cooperative_2020 studied a cooperative ecological driving system for signal corridors, where the CAV switches between a leading vehicle and a following vehicle, and IDM is used to simulate a CAV with network communication failure.
On the road network scale, various scenarios are simulated with IDM as the underlying model. huang_characterizing_2023 considered different signal control schemes and demand loading patterns, and focused on the overall network level to consider the traffic flow patterns brought about by autonomous driving mixed traffic. claes_decentralized_2011 proposed a distributed vehicle path planning method based on predictive information.
@@ 4.5.1 Applications in vehicle communications
In terms of in-vehicle networks, IDM is used as a fundamental model for vehicle operation. wangOptimalFeedbackControlLaw2023a and wang_analytical_2024 have significantly advanced the understanding of cyberattacks on automated vehicles by developing an optimal feedback control law and a comprehensive framework for modeling attacks on adaptive cruise control vehicles. dongDynamicManagerSelectionAssisted2022a introduced a resource allocation algorithm for 5G-V2X platoons, optimizing communication performance under finite block length constraints, crucial for ensuring ultra-reliable low-latency communications. jiaDisturbanceAdaptiveDesignVANETenabledVehicle2014a and jiaNetworkConnectivityPlatoonBasedVehicular2014a delved into VANET-enabled platoon architectures, emphasizing disturbance adaptation and network connectivity, which are vital for effective safety message transmission under varying system parameters. Moreover, segataAutomaticEmergencyBrakingRealistic2013a assessed the integration of automatic emergency braking systems within VANETs, while tianSelfOrganizedRelaySelectionCooperative2017a proposed a decentralized relay selection algorithm based on non-cooperative game theory to enhance cooperative dynamics.
@@ 4.5.2 Applications in civil engineering
ge_intelligent_2022 introduced an advanced traffic flow load monitoring system that integrates machine vision and weigh-in-motion systems, leveraging deep learning for precise vehicle detection and extraction of key parameters for traffic flow load simulation. zhou_hybrid_2024 proposed a two-step Hybrid Virtual-Real Traffic Simulation method to replicate the spatiotemporal distribution of bridge loads, demonstrating the broad applicability of IDM in various interdisciplinary fields.
@@ 5.1 Combination of Big Data
Current IDM research faces several limitations. Despite the widespread use of open-source datasets like the NGSIM, Naples data set (punzoAnalysisComparisonMicroscopicTraffic2005), Hefei data set (jiangExperimentalFeaturesCarfollowingBehavior2015), and the HighD data set (krajewskiHighDDatasetDroneDataset2018), data scarcity remains a significant challenge, affecting both empirical and experimental data acquisition. The limited availability and scope of these datasets make it difficult to extract universal behavioral or macro-level characteristics. Additionally, insufficient data volume constrains model validation and refinement. This situation not only hampers the development of mechanism-driven models like IDM but also limits the effective training of data-driven models, such as those based on machine learning and deep learning. Issues of data quality and diversity are also prominent, as existing datasets may lack representation of specific traffic conditions or regions. Furthermore, inadequate data collection frequency and precision hinder the ability to capture dynamic traffic changes.
Integrating big data into IDM research can significantly address current limitations through both empirical and experimental data approaches. Empirical data, gathered from traffic monitoring systems, floating car data, and intelligent infrastructure, offers a vast and dynamic dataset. This abundance of data provides a robust foundation for testing and validating models across diverse scenarios. By leveraging these large-scale datasets, IDM can be further refined and integrated with machine learning and deep learning techniques, enhancing traffic prediction capabilities. This integration enables more accurate and adaptive models that can efficiently respond to real-time traffic conditions. On the experimental data front, the advantages are equally compelling. Conducting controlled traffic experiments enables researchers to delve deeper into the mechanisms of driver behavior and traffic flow. Experimental data provides the opportunity to isolate and study specific variables, minimizing interference and uncovering fundamental characteristics and mechanisms (tianCarFollowingDynamicsExperiments2023). This detailed insight is crucial for guiding model development and improving theoretical frameworks. By bridging micro-level decision-making processes with macro-level traffic phenomena, experimental data facilitates a comprehensive understanding of traffic dynamics. This approach not only enriches the theoretical foundation of IDM but also supports the development of models that are both innovative and aligned with real-world complexities.
@@ 5.2 Hybrid Modelling Methodology
Hybrid modeling, which combines data-driven and mechanism-driven approaches, offers significant advantages over purely mechanistic or data-driven models. It leverages the strengths of both methodologies: the adaptability and pattern recognition capabilities of data-driven models, and the theoretical rigor and interpretability of mechanism-driven models like (yuanModelingStochasticMicroscopicTraffic2020; yuanMacroscopicTrafficFlowModeling2021). This synergy can lead to more robust and versatile traffic models, capable of addressing complex, real-world scenarios. However, the integration process faces several limitations.
Firstly, the integration often lacks seamless fusion, resulting in limited adaptability and flexibility in practical applications. Secondly, data-driven models and mechanism-driven models often have different underlying structures and assumptions, which makes combining them smoothly a challenge. This can lead to a lack of cohesion in the hybrid model, where components operate in isolation rather than in a unified manner. Additionally, the calibration and tuning processes for each model type can be complex and may not align well, resulting in inefficiencies or conflicts in parameter settings. Additionally, hybrid models can exhibit inconsistent performance across different traffic scenarios, particularly in data-scarce or anomalous situations, affecting their robustness.Moreover, the lack of standardized metrics can lead to unfair comparisons and inconsistent assessments just as wangComparingHundredsMachineLearning2024 pointed out. Furthermore, The challenge of integrating heterogeneous data sources persists, as effectively combining different data types to enhance model performance remains complex. Lastly, computational complexity and real-time applicability are significant constraints; integrated models may demand substantial computational resources, hindering their use in real-time traffic management. Addressing these challenges requires developing more flexible integration frameworks and efficient algorithms to support the advancement and application of IDM models.